OpenStudy (anonymous):

2x - 4y + 3x2 + 7y - 16x

OpenStudy (anonymous):

-8x 3y

OpenStudy (strawberry17):

subtract and add like terms: 3x^2 = 3x^2 2x - 16x = -14x -4y + 7y = 3y Then write them in a in order

OpenStudy (strawberry17):

do you get that?